OpenStudy (anonymous):

solve abs(2x+6)>= abs(9x-8), answer in interval form

OpenStudy (solomonzelman):

\(\large\color{blue}{ \left| 2x+6 \right|= \left| 9x-8 \right| }\) should be converted to, \(\large\color{blue}{ 2x+6 =\pm( 9x-8 ) }\).

OpenStudy (solomonzelman):

where you would have 2 solutions for x.

OpenStudy (camper4834):

are you sure about that?

OpenStudy (camper4834):

\[|2x+6|=\begin{cases} -2x+6\\ 2x+6 \end{cases}\]

OpenStudy (solomonzelman):

I am fairly sure. Reason. when you say that a^2=b^2 (and this statement you can see when you complete the square) you don't say that +-a=+-b No... you just say a=+-b

OpenStudy (solomonzelman):

here, it is sort of the same.

OpenStudy (camper4834):

ahh that gives me another idea

OpenStudy (camper4834):

\[\sqrt{ 2x+6 } ^{ 2 } \ge\sqrt{ 9x-8 } ^{ 2 } \]

OpenStudy (solomonzelman):

yeah, my sign was off, it should be \(\large\color{blue}{ \ge}\) but you got the idea of my first statement though.

OpenStudy (camper4834):

nevermind im not sure how to get plus minus from this

OpenStudy (solomonzelman):

well, you know that if \(\large\color{blue}{ \left| x \right|=a}\), then \(\large\color{blue}{ x=\pm a}\)

OpenStudy (camper4834):

mm so i solved your \(\large\color{blue}{ 2x+6 =\pm( 9x-8 ) }\)

OpenStudy (camper4834):

but it doesn't quite fit with the graph

OpenStudy (camper4834):

neither does the answers at the critical points match

OpenStudy (camper4834):

wait wait

OpenStudy (camper4834):

nevermind it worked XD did bad math

OpenStudy (camper4834):

im curious though about your example a^2=b^2 becoming \(a= \pm b\)